OpenStudy (anonymous):

A cart is moving horizontally along a straight line with constant speed 30m/s. A projectile is to be fired from the moving cart in such a way that it will return to the cart after the cart has moved 80m . At what speed (relative to cartt) must the projectile launched? g=10

OpenStudy (anonymous):

Case 1 : cart 30 * t = 80 t = 8/3s

OpenStudy (anonymous):

Case 2 ; projectile T = 2 * u * sinx/g 8/3 = 2*u *sinx/10

OpenStudy (anonymous):

i'm not really sure...but this is what i got so far: first, we know Range R = 80m and Time of flight T = 8/3 sec. let u be the initial speed. so using their respective formulae, i got ucosx = R/T ucosx = 30.................since R/T = 30 u = 30/cosx the speed of projectile relative to that of cart will be (speed of cart) - (speed of projectile)...........right? = 30 - 30/cosx = 30(1 - secx) is this right?

OpenStudy (anonymous):

The options are 10 10sqrt8 40/3 none of these

OpenStudy (anonymous):

@Vaidehi09

OpenStudy (anonymous):

lol..yes, i know. we still need to find x.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hw do we find that

OpenStudy (anonymous):

that's what i'm trying to work out.

OpenStudy (anonymous):

@waterineyes is there a flaw in my approach? if yes, then can u please point it out?

OpenStudy (dominusscholae):

Wait, we don't necessarily have to find x. Assuming from the problem, you're firing the ball straight up while moving on the cart. What we should be focusing on is the vertical velocity, since that is the speed RELATIVE TO THE CART. We want to find how it takes for the ball to reach its maximum height. We use the formula:\[Vy - Gt = 0\], G = 10, to do this. Since we are talking about the maximum height, that means the ball has take half of its time to fall back into the cart. Thus, our time should be 4/3. Doing so yields \[Vy = 40/3\]. Thus The answer is 40/3.

OpenStudy (anonymous):

i dont think that is correct if he throws vertically upwards how the projectile will cover 80m

OpenStudy (anonymous):

range will be 0 then...

OpenStudy (anonymous):

yes, there should be an angle of projection.

OpenStudy (dominusscholae):

Perhaps the question is unclear to this, but it seems that the cart, because the cart is moving, will move with 30 m/s as well.

OpenStudy (dominusscholae):

It does say the cart is moving.

OpenStudy (anonymous):

so the projectile will have intial speed = speed of cart + muzzzle power

OpenStudy (anonymous):

if we go by what @dominusscholae says, then the eq should be Vy sinx = gt so, Vy sinx = 40/3

OpenStudy (anonymous):

initial horizontal velocity of the projectile will be equal to 30ms^-1?

OpenStudy (dominusscholae):

Yes. The problem is that it doesn't say whether you're shooting upward. You have to make the assumption that it is, and I made it on the basis of your work. The speed in the x-direction, though, would only come from the cart, while the speed in the y-direction comes from the muzzle shot.

OpenStudy (anonymous):

we need to find the angle of projection!

OpenStudy (anonymous):

i think the answer is 'none of these'....

OpenStudy (dominusscholae):

I don't know. Is that all the question gives? Far as I know, there's a ton of interpretations and right answers that relate to this question if there is. What is confusing is that in case 1, you didn't factor in a horizontal velocity. Thus I was led to assume we were shooting it up.

OpenStudy (anonymous):

case 1 is for the cart. it has velocity only in one direction - horizontal.

OpenStudy (anonymous):

and the horizontal velocity of the projectile will be 30ms as well?

OpenStudy (anonymous):

i don't think so.

OpenStudy (dominusscholae):

It would have a horizontal velocity of 30 ms. You also have to consider that the problem asked for the speed relative to the cart, which means that the only velocity to find is the vertical velocity of the projectile.

OpenStudy (anonymous):